# MobileRL: Advancing Mobile Use Agents with Adaptive Online Reinforcement Learning

### v1

@@ 2.2.1 Shortest-Path Reward Adjustment (SPA)
$R^{\mathrm{SPA}}(s_{t},a_{t})=\begin{cases}\displaystyle 1-\alpha\,\frac{T_{i}-T_{\min}}{T_{i}},&r(\tau_{i})=1,\\[6.0pt] 0,&r(\tau_{i})=0,\end{cases}\qquad T_{\min}=\min_{\tau_{j}\in\mathcal{T}_{\mathrm{succ}}}|\tau_{j}|,\;\alpha\in(0,1],$ (1)
@@
Trajectory-level Advantage.
Given a task, we sample $G$ trajectories $\mathcal{G}={\tau_{1},\ldots,\tau_{G}}$. For trajectory $\tau_{i}$ with $T_{i}$ valid steps, we define the trajectory-level advantage as the average of standardized SPA rewards:
$A_{\mathrm{traj}}(\tau_{i})=\frac{1}{T_{i}}\sum_{t=1}^{T_{i}}\frac{R^{\mathrm{SPA}}(s_{i,t},a_{i,t})-\mathrm{avg}{j\in\mathcal{G}},R^{\mathrm{SPA}}(s{j,t},a_{j,t})}{\mathrm{std}{j\in\mathcal{G}},R^{\mathrm{SPA}}(s{j,t},a_{j,t})}.$ (2)
@@ 2.2.2 Difficulty–Adaptive Positive Replay (AdaPR)
Difficulty-Adaptive Positive Replay (AdaPR) strategically retains and reuses challenging, high-value trajectories while blending them with fresh on-policy samples. In sparse-reward mobile environments, successful but difficult rollouts are rare yet highly informative; leveraging them effectively strengthens the learning signal and stabilizes policy updates. We introduce three key components of AdaPR: buffer construction for high-quality trajectory selection, mixture sampling to balance replay and exploration, and negative rollout pruning to reduce noise.
@@
At iteration $t$, the rollout set is $\mathcal{T}_{t}=\{\tau^{(1)},\dots,\tau^{(N)}\}$, collected under the current policy $\pi_{\theta_{t}}$. We compute $A_{\mathrm{traj}}(\tau)$ via equation 2 and insert the top $\kappa$ trajectories into the replay buffer $\mathcal{B}$.
@@
Each policy update is performed on a mini-batch of $M$ trajectories obtained from the mixture distribution
$q(\tau)\;=\;\gamma\,p_{\mathcal{B}}(\tau)\;+\;(1-\gamma)\,p_{\mathrm{on}}(\tau)$ (3)
where $p_{\mathrm{on}}$ is the on-policy distribution $\pi_{\theta_{t}}$ and $p_{\mathcal{B}}$ is the empirical distribution over $\mathcal{B}$. To keep the replay contribution under control, at most $\gamma M$ trajectories with the highest current advantage are drawn from $\mathcal{B}$, preserving on-policy diversity.
Pruning Negative Rollouts.
@@ 2.2.3 Failure Curriculum Filtering (FCF)
To avoid repeatedly sampling tasks that yield zero reward—which wastes computation and hinders the collection of positive advantage data, we propose failure curriculum filtering. In FCF, any task producing all-zero rewards for two consecutive epochs enters a three-epoch cooldown, during which its sampling probability is reduced according to $w_{\text{task}}=\exp(-f)$, where $f$ is the number of consecutive failure epochs. Tasks with consistently high failure counts are permanently removed from the sampling pool. For stability, failure histories from previous training are retained.
@@
In summary, MobileRL consists of reasoning-free SFT, reasoning-augmented warm-up, and difficulty-adaptive RL for training mobile GUI agents. Reasoning-free SFT helps build a strong action foundation from expert demonstrations, while reasoning SFT adds intermediate reasoning to improve instruction following and policy transparency. On top of this initialization, agentic RL with AdaGRPO addresses the challenges of sparse terminal rewards, heavy-tailed task difficulty, and expensive sampling. Specifically, SPA reshapes terminal rewards for denser feedback, AdaPR strategically reuses challenging successful trajectories, and FCF filters out persistently-unsolvable tasks.
@@ 3 Experiments
3.1 Experiments Settings
Datasets and Benchmarks.
We evaluate on two interactive Android benchmarks: AndroidWorld with 116 tasks across 20 apps (Rawles et al., 2024) and AndroidLab with 138 tasks across 9 apps (Xu et al., 2024), totaling 254 tasks that span diverse, real-world goals. Both sets offer instruction-following environments where agents act to complete tasks. We use AndroidWorld’s rule-based trajectory rewards; for AndroidLab, which lacks training rewards, we generate signals via a VLM-based reward model (see Appendix C).
Baselines.
Our baselines encompass both closed- and open-source agents and models. The closed-source LLMs include GPT-4o-2024-11-20 (OpenAI, 2023) and Claude-Sonnet-4-20250514-thinking (Anthropic, 2023), and closed-source agents UI-Tars-1.5 (Qin et al., 2025) and AutoGLM (Liu et al., 2024). The open-source VLMs, including Qwen2.5-VL-7B-Instruct (Bai et al., 2025), GLM-4.1V-9B-Thinking (Team et al., 2025), UI-Tars-7B (Qin et al., 2025), V-Droid (Dai et al., 2025) and UI-Genie-Agent (Xiao et al., 2025), are used.
3.2 Main Results
@@
w/ GLM-4.1V-9B-Base 9B 75.8 46.8
We evaluate MobileRL by using Qwen2.5-VL-7B and GLM-4.1V-9B-Base as backbones on AndroidWorld and AndroidLab. As shown in Table 1, MobileRL significantly outperforms both proprietary models (e.g., GPT-4o and Claude-Sonnet-4) and open-source models (e.g., V-Droid and UI-Tars-7B). With Qwen2.5-VL-7B as the backbone, MobileRL achieves 72.0%on AndroidWorld and 42.5%on AndroidLab, outperforming previous state-of-the-art methods. By using GLM-4.1V-9B as the backbone, the performance of MobileRL further improves to 75.8%on AndroidWorld and 46.8%on AndroidLab, achieving the best results across all models.
@@ 3.3 Ablation Study
   + Reasoning-Free SFT 48.9+41.2 39.8+29.7
    + Reasoning SFT 66.1+17.2 40.3+0.5
     + AdaGRPO (MobileRL-9B) 75.8+9.7 46.8+6.5
@@
Figure 3: Ablation studies of the MobileRL framework and its AdaGRPO algorithm. All test set results are averaged over three runs to mitigate the impact of randomness.
To evaluate the contributions of the MobileRL framework and the components of the AdaGRPO algorithm, we start with two base models and progressively apply Reasoning-Free SFT, Reasoning SFT, and AdaGRPO. Then, using the Qwen2.5-VL-7B-Instruct model trained with Reasoning-Free SFT and Reasoning SFT as the initialization point, we conduct an analysis of the impact of each component of AdaGRPO—AdaPR, SPA, and FCF.
@@
We summarize stage-wise gains in success rate (SR) in Table 3(a). For Qwen2.5-VL-7B, the combined improvements of MobileRL are +44.4% on AndroidWorld and +32.4% on AndroidLab; For GLM-4.1V-9B, the overall gains are +68.1% and +36.7%. Overall, Reasoning-Free SFT delivers a strong initial lift and Reasoning SFT offers additional improvements. Building upon the strong foundation established by the preceding stages, the AdaGRPO stage further augments the final performance, achieving an improvement exceeding 10% on the AndroidWorld dataset, where rewards are verifiable, and over 5% on the AndroidLab test set, which is evaluated using the reward model.
@@
We report on-policy trajectory reward curves during training (excluding replayed trajectories) in Figure 3(b) and the final success rates on the AndroidWorld test set in Table 3(c). To avoid bias from the AndroidLab reward model, we use only AndroidWorld in this study.
Each component of AdaGRPO contributes to improving the performance of MobileRL. Specifically, we have the following observations:
@@
  • Effect of SPA. Noticeable gains of SPA appear after roughly 60 steps, likely because the lack of step-wise control leads to overly long trajectories. Section 3.3 confirms that SPA effectively constrains trajectory length.
Figure 4: Effect of reasoning-free SFT evaluated on AndroidWorld.
@@
We apply supervised fine-tuning on the expert dataset without reasoning traces, which we term Reasoning-Free SFT. The fine-tuning data in this stage contains only action sequences, without intermediate thought processes. This raises a key question: Is fine-tuning with expert data that lacks reasoning still beneficial?
We compare MobileRL (Reasoning-Free SFT + Reasoning SFT + AdaGRPO) with MobileRL without Reasoning-Free SFT in Figure 5. Interestingly, our experiments show that incorporating Reasoning-Free SFT consistently improves performance on AndroidWorld. It suggests that without explicit reasoning traces, (Reasoning-Free) SFT contributes to stabilizing training and enhancing final results for MobileRL.
Effect of Pruning Negative Trajectories.
@@
We divide the AndroidWorld test set by rounded-up Complexity (Rawles et al., 2024): complexity level 1 (1–10 steps), complexity level 2 (11–20), complexity level 3 (21–30), and complexity level 4+ (>>30). We run eight test trials at temperature $1.0$ and report pass@1/2/4/8 in Figure 6.
Our method yields consistent gains at all difficulty levels, with improvements increasing alongside complexity. Notably, post-RL pass@1 exceeds pre-RL pass@8, indicating substantial effectiveness. The relatively modest changes in pass@8 suggest exploration is preserved but better targeted. Consistent with AdaPR ’s design for heavy-tailed difficulty, the largest gains occur on high-complexity tasks (complexity levels 3 and 4).
@@
Figure 7: Win rates of MobileRL vs. MobileRL w/o SPA, where a win means completing a task with fewer steps. $n$ denotes the number of task templates per category. Categories: All (all templates); C1, C2, C3, C4 (complexity levels 1,2,3,4, respectively); BC/BW (both correct/both wrong); Others (exactly one method correct).
@@
Although SPA has the smallest impact on overall accuracy in the ablation, its effect on step efficiency is clear. As shown in Figure 7, partitioning tasks by complexity reveals that MobileRL with SPA consistently completes tasks in fewer steps across all difficulty levels. Moreover, when we compare cases where both models are correct (BC), both are wrong (BW), and those where only one is correct, MobileRL with SPA more frequently yields shorter trajectories in every group—most prominently in the BW setting—indicating that SPA systematically reduces completion length regardless of final correctness.
@@ 4 Related Work
Recent work leverages powerful language models to build agents that operate real PCs and phones (Agashe et al., 2025; Qin et al., 2025; Lai et al., 2025b), including Android agents that perceive GUIs and act via taps, swipes, and text (Toyama et al., 2021; Xu et al., 2024). To improve action prediction and learning, frameworks explore multimodal exploration (Yang et al., 2023), modular reasoning (Lai et al., 2025a), verifier-driven control (Dai et al., 2025), and small-LM code-based execution (Wen et al., 2025). Yet many systems still rely on offline RL or single-turn data: DigiRL uses offline demonstrations (Bai et al., 2024); U1-R1 trains on single-step episodes (Lu et al., 2025); and UI-Tars applies DPO in an offline regime (Qin et al., 2025), leaving online, multi-turn RL for adaptive mobile agents largely unexplored.
@@
Benchmarking generally follows two tracks. Static or replay-based settings—AndroidControl, Android in the Wild, MobileAgentBench, and Mobile-Bench—offer many tasks but lack real-time interaction (Li et al., 2024; Rawles et al., 2023; Wang et al., 2024; Deng et al., 2024). Interactive emulator environments—AndroidWorld, AndroidLab, and B-MOCA—span diverse apps and realistic tasks yet remain challenging for current agents (Rawles et al., 2024; Xu et al., 2024; Lee et al., 2024). To the best of our knowledge, all public mobile GUI benchmarks to date target the Android operating system.
@@ 5 Conclusion
In this work, we present MobileRL, an agentic RL framework that advances mobile GUI agents. It achieves this by combining staged initialization with an adaptive reinforcement learning algorithm (AdaGRPO). Training begins with reasoning-free SFT on large-scale expert demonstrations to establish a solid task foundation, followed by a reasoning-augmented warm-up stage that adds intermediate rationales to improve transparency and reduce cold-start exploration costs. Building on this, we introduce Difficulty–Adaptive GRPO (AdaGRPO), which enhances GRPO with shortest-path reward adjustment, adaptive positive replay, and failure curriculum filtering. These three strategies improve sample efficiency and guide policies toward more accurate and efficient task completion. Experiments on AndroidWorld and AndroidLab demonstrate that MobileRL with open models significantly outperforms both open-source and proprietary baselines.
For future work, we aim to improve the broader generalization of MobileRL across diverse mobile platforms, application domains, and unseen UI layouts. In addition, we plan to develop a general verification model that can automatically assess the correctness, reliability, and safety of agent actions, enabling more robust and trustworthy deployment in real-world scenarios.

### v2

@@ 2.2.1 Shortest‐Path Reward Adjustment (SPA)
$R^{\mathrm{SPA}}(s_{t},a_{t})=r(\tau_{i})\left(1-\alpha\,\frac{T_{i}-T_{\min}}{T_{i}}\right),\qquad T_{\min}=\min_{\tau_{j}\in\mathcal{T}_{\mathrm{succ}}}|\tau_{j}|,\;\alpha\in(0,1].$ (2)
@@ 2.2.2 Difficulty–Adaptive Positive Replay (AdaPR)
In sparse-reward mobile environments, successful yet challenging rollouts are rare but highly informative; effectively leveraging them enhances the learning signal and stabilizes policy updates. Therefore, inspired by the paradigm of experience replay (Mnih et al., 2015; Zha et al., 2019), we introduce difficulty-adaptive positive replay (AdaPR) strategically to retain and reuse challenging, high-value trajectories while blending them with fresh on-policy samples. We introduce key components of AdaPR: buffer construction for high-quality trajectory selection and mixture sampling to balance replay and exploration.
@@
At iteration $t$, the rollout set is $\mathcal{T}_{t}=\{\tau^{(1)},\dots,\tau^{(N)}\}$, collected under the current policy $\pi_{\theta_{t}}$. We compute the trajectory-level advantage and insert the top $\kappa$ trajectories into the replay buffer $\mathcal{B}$.
@@
Each policy update is performed on a mini‑batch of $M$ trajectories obtained from the mixture distribution $q(\tau)\;=\;\gamma\,p_{\mathcal{B}}(\tau)\;+\;(1-\gamma)\,p_{\mathrm{on}}(\tau)$ where $p_{\mathrm{on}}$ is the on‑policy distribution $\pi_{\theta_{t}}$ and $p_{\mathcal{B}}$ is the empirical distribution over $\mathcal{B}$. To keep the replay contribution under control, at most $\gamma M$ trajectories with the highest current advantage are drawn from $\mathcal{B}$, preserving on‑policy diversity.
@@ 2.2.3 Failure Curriculum Filtering (FCF)
To avoid repeatedly sampling tasks that yield zero reward—which wastes computation and hinders the collection of positive advantage data, we propose failure curriculum filtering. In FCF, any task producing all-zero rewards for two consecutive epochs enters a three-epoch cooldown, during which its sampling probability is reduced according to $w_{\text{task}}=\exp(-f)$, where $f$ is the number of consecutive failure epochs; after this cooldown period, the task is permanently removed. This method can be regarded as a variant of curriculum sampling (Matiisen et al., 2019; Narvekar et al., 2020). To avoid excessive hyperparameter tuning, we simplify it to target only the exclusion of the most difficult subset of data. Tasks with consistently high failure counts are permanently removed from the sampling pool. For stability, failure histories from previous training are retained.
@@
In summary, MobileRL consists of reasoning-free SFT, reasoning SFT, and difficulty-adaptive RL for training mobile GUI agents. Reasoning-free SFT helps build a strong action foundation from expert demonstrations, while reasoning SFT adds intermediate reasoning to improve instruction following and policy transparency. On top of this initialization, agentic RL with AdaGRPO addresses the challenges of sparse terminal rewards, heavy-tailed task difficulty, and expensive sampling. Specifically, SPA reshapes terminal rewards for denser feedback, AdaPR strategically reuses challenging successful trajectories, and FCF filters out persistently-unsolvable tasks.
@@ 3 Experiments
w/ GLM-4.1V-9B-Base 9B 80.2 53.6
3.1 Experiments Settings
Datasets and Benchmarks.
In the two-stage SFT process, we utilize data from human annotations and AndroidControl (Li et al., 2024), constructing 97.9k and 23.6k training steps, respectively. During the RL stage, we employ interactive training sets from AndroidWorld (Rawles et al., 2024) and AndroidLab (Xu et al., 2024), consisting of 2,000 and 1,103 tasks. These were paired with verifiable rewards and VLM-based reward models (cf. Appendix D.3.2). We evaluate on two interactive Mobile benchmarks: AndroidWorld, which includes 116 tasks across 20 apps, and AndroidLab, which comprises 138 tasks across 9 apps. Both benchmarks provide interactive environments. We further assess performance on the static dataset AndroidControl (Li et al., 2024), which contains 8,444 test samples. Due to space limitations, additional details on the training and test datasets are provided in Appendix D.
Baselines.
Our baselines encompass both closed‑ and open‑source agents and models. The closed‑source LLMs include GPT‑4o‑2024‑11‑20 (OpenAI, 2023) and Claude-Sonnet-4-20250514-thinking (Anthropic, 2023), and closed‑source agents UI‑Tars‑1.5 (Qin et al., 2025) and AutoGLM (Liu et al., 2024). The open‑source VLMs, including Qwen2.5‑VL‑7B-Instruct (Bai et al., 2025), GLM-4.1V-9B-Thinking (Team et al., 2025), UI‑Tars‑7B (Qin et al., 2025), V‑Droid (Dai et al., 2025) and UI-Genie-Agent (Xiao et al., 2025), are used.
3.2 Main Results
We evaluate MobileRL by using Qwen2.5-VL-7B and GLM-4.1V-9B-Base as backbones on online interactive benchmarks, AndroidWorld and AndroidLab. As shown in Table 1, MobileRL significantly outperforms previous results. With Qwen2.5-VL-7B as the backbone, MobileRL achieves 72.0%  on AndroidWorld and 42.5%  on AndroidLab, outperforming previous state-of-the-art methods. By using GLM-4.1V-9B as the backbone, the performance of MobileRL further improves to 80.2%  on AndroidWorld and 53.6%  on AndroidLab, achieving the best results across all models.
@@ 3.3 Ablation Study
   + Reasoning-Free SFT 48.1+40.4 42.7+32.6
    + Reasoning SFT 66.2+18.1 45.0+2.3
     + AdaGRPO (MobileRL-9B) 80.2+14.0 53.6+8.6
@@
Figure 3: Ablation studies of the MobileRL framework and its AdaGRPO algorithm. We use the Reasoning SFT model with Qwen2.5-VL-7B-Instruct backbone for the ablation of the AdaGRPO algorithm. All test set results are averaged over three runs to mitigate the impact of randomness.
To evaluate the contributions of the MobileRL framework and the components of the AdaGRPO algorithm, we start with two base models and progressively apply Reasoning-Free SFT, Reasoning SFT, and AdaGRPO. Then, using the Qwen2.5-VL-7B-Instruct model trained with Reasoning SFT as the initialization point and on the AndroidWorld training set, we conduct an analysis of the impact of each component of AdaGRPO—AdaPR, SPA, and FCF.
@@
We summarize stage-wise gains in success rate (SR) in Table 3(a). For Qwen2.5-VL-7B, the combined improvements of MobileRL are +44.4% on AndroidWorld and +32.4% on AndroidLab; For GLM-4.1V-9B, the overall gains are +72.5% and +43.5%. Overall, Reasoning-Free SFT delivers a strong initial lift, and Reasoning SFT offers additional improvements. Building upon the strong foundation established by the preceding stages, the AdaGRPO stage further augments the final performance, achieving an improvement exceeding 10% on the AndroidWorld dataset and over 5% on the AndroidLab test set.
@@
We report on-policy trajectory reward curves during training (excluding replayed trajectories) in Figure 3(b) and the success rates on the AndroidWorld in Table 3(c). To avoid bias from the AndroidLab reward model, we use only AndroidWorld in this study. Each component of AdaGRPO contributes to improving the performance of MobileRL. Specifically, we have the following observations:
@@
  • Effect of SPA. Noticeable gains of SPA appear after roughly 60 steps, likely because the lack of step-wise control leads to overly long trajectories.
@@
We apply supervised fine-tuning on the expert dataset without reasoning traces, which we term Reasoning-Free SFT. This raises a key question: Is fine-tuning with expert data that lacks reasoning still beneficial?
We compare MobileRL (Reasoning-Free SFT + Reasoning SFT + AdaGRPO) with MobileRL without Reasoning-Free SFT in Figure 3. Interestingly, our experiments show that incorporating Reasoning-Free SFT consistently improves performance on AndroidWorld. It suggests that without explicit reasoning traces, (Reasoning-Free) SFT contributes to enhancing final results for MobileRL.
Table 2: Results on the AndroidControl (Li et al., 2024). Note that data from AndroidControl was not included during the RL stage.
Model AndroidControl-Low AndroidControl-High
Type Grounding SR Type Grounding SR
UI-Tars-7B 98.0 89.3 90.8 83.7 80.5 72.5
UI-Tars-72B 98.1 90.8 91.3 85.2 81.5 74.7
UI-Genie-7B 98.1 89.3 94.3 83.5 82.9 74.5
UI-Genie-72B 98.3 95.4 94.8 84.9 86.3 77.0
MobileRL-7B
   w/o RL 96.7 93.7 91.9 87.0 73.6 72.8
   w/ RL 96.9 92.4 91.4 86.3 71.5 71.3
MobileRL-9B
   w/o RL 98.6 95.8 95.9 89.8 78.1 77.5
   w/ RL 97.9 94.6 94.3 87.0 74.5 73.9
Table 3: Effect of reasoning-free SFT evaluated on AndroidWorld.
Performance on Offline Dataset.
As shown in Appendix D, Reasoning-Free and Reasoning SFT stages include training data from AndroidControl, which has been converted into MobileRL format. In Table 3, we present the results of MobileRL on the AndroidControl test set. The SR score of MobileRL-9B w/o RL version surpasses all previous models, achieving the state of the art, while the RL version largely maintains the score.
@@
We divide the AndroidWorld test set by rounded-up Complexity (Rawles et al., 2024): complexity level 1 ($\leq$10 steps), level 2 (11–20), level 3 (21–30), and level 4+ (>>30). We run eight test trials at temperature $1.0$ and report pass@1/2/4/8 in Figure 5. MobileRL yields consistent gains at all complexity levels. Notably, post-RL pass@1 exceeds pre-RL pass@8, indicating substantial effectiveness. Consistent with AdaPR’s design for heavy-tailed difficulty, larger gains occur on high-complexity tasks.
@@
Figure 5: Win rate of MobileRL vs. MobileRL w/o SPA, where a win means completing a task with fewer steps. $n$ denotes the number of task templates per category. Categories: All (all templates); C1–C4 (complexity levels 1–4); BC/BW (both methods correct/wrong); Others (exactly one method correct).
@@
Although SPA has the smallest impact on overall accuracy in the ablation, its effect on step efficiency is clear. As shown in Figure 5, partitioning tasks by complexity reveals that MobileRL with SPA consistently completes tasks in fewer steps across all difficulty levels. Moreover, when we compare cases where both models are correct (BC), both are wrong (BW), and those where only one is correct, MobileRL with SPA more frequently yields shorter trajectories in every group.
@@ 4 Related Work
Recent work leverages powerful language models to build agents that operate real PCs and phones (Agashe et al., 2025; Qin et al., 2025; Lai et al., 2025b), including Android agents that perceive GUIs and act via taps, swipes, and text (Toyama et al., 2021; Xu et al., 2024). To improve action prediction and learning, frameworks explore multimodal exploration (Yang et al., 2023), modular reasoning (Lai et al., 2025a), verifier-driven control (Dai et al., 2025), and small-LM code-based execution (Wen et al., 2025). Yet many systems still rely on offline RL or single-turn data: DigiRL uses offline demonstrations (Bai et al., 2024); U1-R1 trains on single-step episodes (Lu et al., 2025); and UI-Tars applies DPO in an offline regime (Qin et al., 2025), leaving online, multi-turn RL for adaptive mobile agents unexplored. Studies emphasize realistic applications: AppAgent (Yang et al., 2023) evaluates closed-source models on real-world apps, while A3 (Chai et al., 2025) offers a realistic app suite with an autonomous evaluation protocol that reduces human effort.
@@
Benchmarking generally follows two tracks. Static or replay-based settings—AndroidControl (Li et al., 2024), Android in the Wild (Rawles et al., 2023), MobileAgentBench (Wang et al., 2024), and Mobile-Bench (Deng et al., 2024) offer plenty of tasks and trajectories. These benchmarks fall short for real-world evaluation, as fixed trajectories and screens limit agents’ ability to handle uncertainty and exploration. Interactive emulator environments—AndroidWorld (Rawles et al., 2024), AndroidLab (Xu et al., 2024), and B-MOCA (Lee et al., 2024)—span diverse apps and realistic tasks, yet remain challenging for current agents. Current mobile GUI benchmarks are limited: they lack asynchronous, parallel VM interaction for scalable training. To the best of our knowledge, all public mobile GUI benchmarks to date target the Android operating system.
@@ 5 Conclusion
In this work, we present MobileRL, an agentic RL framework that advances mobile GUI agents. It achieves this by combining staged initialization with an adaptive reinforcement learning algorithm (AdaGRPO). Training begins with reasoning-free SFT on large-scale expert demonstrations, followed by a reasoning SFT stage that adds intermediate rationales and reduces cold-start exploration costs. Building on this, we introduce Difficulty–ADAptive GRPO (AdaGRPO), which enhances GRPO with shortest-path reward adjustment, adaptive positive replay, and failure curriculum filtering. These three strategies improve sample efficiency and guide policies toward more accurate and efficient task completion. Experiments on AndroidWorld and AndroidLab demonstrate that MobileRL with open models significantly outperforms both open-source and proprietary baselines.